OpenStudy (anonymous):

Explain how you would find the exact value of:

OpenStudy (anonymous):

\[\sqrt{2} \over \sqrt{8}\]

Parth (parthkohli):

Hmm.... \(\Large \color{Black}{\Rightarrow {\sqrt{2} \times \sqrt{8} \over 8}}\) \(\Large \color{Black}{\Rightarrow {\sqrt{16} \over8}}\) \(\Large \color{Black}{\Rightarrow {4 \over 8} }\) \(\Large \color{Black}{\Rightarrow {1 \over 2} = 0.5 = 50\% = {5 \over 10} }\)

Parth (parthkohli):

1/2, 0.5, 50% and 5/10 are all equal

OpenStudy (anonymous):

alright thanks for all the help parth